how do you find f(x)=ax+b in which f(2)=1 and f(-3(=11?
from mathematics for the international student:Mathematics SL-section 1c

Answer by alka001(14)

f(x)=ax+b
put x=2
f(2)=a*2+b=1
=2a+b=1 ...............(1)
put x=-3
f(-3)=a*-3+b=11
=-3a+b=11 .............(2)
(1)-(2)...5a=-10
a=-2
put a=-2in (1)
2*-2+b=1
-4+b=1
b=1+4
b=5
so,f(x)=-2a+5